Dear,
Here is a summary of CREAM-IV flight preparation status on the 18th of November at McMurdo.
- In the weather briefing, nothing much changed yet. We still have a windy weather expected to coming on Thursday (Wed. in US) when we plan to do the End-to-End test with the ROCC/ESC/SOC. Will prepare the test as planned and change it if weather becomes bad.
- The pattern of stratosphere is currently about 2 days later than 2007 but still expect to have the vortex formed around 10-15th of December.
- The instrument CPT is complete. The calorimeter, t/b SCD, Chercam, TCD tests are done today. Everything works fine except the following items.
-
- TCD MXL PM#1 had a unstable TCD channel and the TCD board replaced. But after that, the ADC data are not shown. Will be replaced and tested whenever time is available after the LPT with the CDM.
- The top SCD data is unstable in the gain runs as we have seen before. Will be checked before the CPT with the CDM starts next Wednesday.
- The calorimeter MB#7 temperature HSK display is incorrect as we have been seen. Will replace the board before the next CPT also.
- The replaced battery for the CMD had a problem with thermostat. The thermostat is replaced and tested OK.
- The CDM is mounted under the pallet.
- The building UPS is available from around noon.
- Tomorrow, the CMD will be connected and a LPT with the CDM will be done.



